# A Survey on Data Selection for LLM Instruction Tuning

### v1

@@ Abstract
Instruction tuning is a vital step of training large language models (LLM), so how to enhance the effect of instruction tuning has received increased attention. Existing works indicate that the quality of the dataset is more crucial than the quantity during instruction tuning of LLM. Therefore, recently a lot of studies focus on exploring the methods of selecting high-quality subset from instruction datasets, aiming to reduce training costs and enhance the instruction-following capabilities of LLMs. This paper presents a comprehensive survey on data selection for LLM instruction tuning. Firstly, we introduce the wildly used instruction datasets. Then, we propose a new taxonomy of the data selection methods and provide a detailed introduction of recent advances,and the evaluation strategies and results of data selection methods are also elaborated in detail. Finally, we emphasize the open challenges and present new frontiers of this task.
1 Introduction
Large language models (e.g. PaLMChowdhery et al. (2023), GPT-4 OpenAI (2023), and LLaMa Touvron et al. (2023)) exhibit remarkable proficiency across a broad spectrum of language understanding and generation tasks, adhering to human instructions with effectiveness and safety. During training process, LLMs typically involves two fundamental steps Ouyang et al. (2022a): pretraining on the large-scale corpus, fine-tuning on instruction datasets. In these steps, fine-tuning on instruction datasets, also referred to instruction tuning, plays a crucial role on aligning LLMs with human instruction. Through training LLMs on (instruction, output) datasets, instruction tuning effectively bridges the gap between LLMs and various of human intents Zhang et al. (2023). Specifically, instruction tuning could align the output of LLMs to human preferences, which enhances the controllability and safety of LLMs. Another benefit is instruction tuning can enable large models to adapt more quickly to specific domains or learn specialized knowledge without requiring significant computational resources and architectural changes.
ln the early researches, the work of instruction tuning mainly focus on construct large-scale instrcution dataset, and creating instruction datasets can be done in two primary ways. One is transforming text-label pairs from existing annotated natural language datasets to instruction-output pairs through templates, such as P3Sanh et al. (2022). Another method involves using LLMs like GPT-3.5-Turbo to generate outputs for given instructions, such as self-instructWang et al. (2023). Although there are various large-scale instruction dataset created by kinds of methods,they often have limitations in terms of quantity, diversity, and creativity. Additionally, enhancing the ability to follow instructions and handling unexpected responses on a large scale of instruction datasets is a current issue that remains to be resolved.
Therefore, selecting the appropriate dataset is crucial for the instruction fine-tuning phase. While instruction fine-tuning primarily relies on a large volume of data, research such as LIMAZhou et al. (2023a) indicates that data quality is more critical than quantity. They demonstrated a significant performance improvement in LLMs using only 1k high-quality instruction data. This finding suggests that LLMs have already acquired world knowledge during the pre-training phase, and the instruction tuning stage requires only a small amount of high-quality instruction data to generate high-quality responses.
Manual instruction data selection often involves high costs and introduces human bias. As a result, creating automated methods for efficiently selecting instruction data has become crucial. However, this task is challenging due to the complex factors and multidimensional considerations involved. For example, it is difficult to assess the quality of individual instructions and ensure the overall diversity of the selected data. Another challenge is reducing costs and improving the efficiency of the selection process. In light of these factors, various data selection methods have been developed.Some methods use a system of indicators to evaluate individual data points, while others rely on trainable LLMs or powerful external LLMs. These exploit the LLMs’ own capabilities to select instructions. Additionally, methods that use smaller models and design a comprehensive process to achieve a balanced effect across all aspects are noteworthy. These methods have shown promising results. For instance, the IFDLi et al. (2023a) method significantly outperforms the Alpaca model by using only about 5% of the Alpaca dataset, and it also surpasses the WizardLM model by about 10%. Using high-quality subsets for fine-tuning not only enhances the instruction-following skills of LLMs but also significantly reduces computational costs and time.
@@ 1 Introduction
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6em,font=,, where level=2text width=6em,font=,, where level=3text width=8em,font=,, where level=4text width=5em,font=,, [ Data Selection for Instruction Tuning, ver [Instruction
Sets (§2) [ Alpaca Taori et al. (2023), WizardLM Xu et al. (2023a),
Dolly-v2 Conover et al. (2023), P3 Sanh et al. (2022),
LIMA Zhou et al. (2023a), self-instruct Wang et al. (2023) , leaf, text width=37em ] ] [Data Selection
Methods (§3) [System of
indicators
3.1) [ INSTRUCTMINING Cao et al. (2023), InstructionGPT-4 Wei et al. (2023),
@@
DQ Zhou et al. (2023b), , leaf, text width=37em ] ] [Trainable
LLMs (§3.2) [ IFD Li et al. (2023a), Instruction Backtranslation Li et al. (2023b),
Nuggets Li et al. (2023c), DIVERSEEVOL Wu et al. (2023),
@@
TEGIT Chen et al. (2023c), ActiveIT Kung et al. (2023) , leaf, text width=37em ] ] [Powerful
LLMs (§3.3) [ AlpaGasus Chen et al. (2023b), INSTAG Lu et al. (2023), LIFT Xu et al. (2023b),
DEITA Liu et al. (2023), tree-instruct Zhao et al. (2023), WaveCoder Yu et al. (2023) , leaf, text width=37em ] ] [Small
@@
Models (§3.4) [ MoDS Du et al. (2023), Coreset-based Selection Chen et al. (2023a), , leaf, text width=37em ] ] ] [Evaluation
Methods (§4) [ Wining rate (§4.1), Inner Comparison (§4.2), External Comparison (§4.3), leaf, text width=37em ] ] ] ]
Figure 1: Overview of Data Selection for LLM Instruction Tuning.
@@
This paper provides a comprehensive review of existing works on data selection methods for the instruction tuning of LLM. To facilities the community, we maintain a paper list [^1] https://github.com/Bolin97/awesome-instruction-selector, collect commonly instruction sets for data selection. Section 2 describes the mainstream datasets with different source and construction methods used in instruction tuning and Section 3 describes four types data selection methods in details: indicators sets, trainable LLMs, powerful LLMs and small models. Section 4 presents the evaluation methods and shows the results of different instruction selection methods. Section 5 summarizes the paper and emphasize the open challenges and future direction of the instruction selection.
2 Instruction Sets
@@
Various instruction tuning datasets (e.g. Self-Instruct and Alpaca), generated by LLMs, offer a wealth of samples without human labor, but their data quality depends on the performance of LLMs and is uncertain. Conversely, the manually curated datasets (e.g. LIMA and Dolly) obtain higher quality through meticulous human selection but are potentially influenced by human biases. Alternative dataset construction methods, like prompt mapping and evol-instruct, aim to enhance dataset quality and diversity but introduce new challenges in quality assurance. This variability in dataset construction and sourcing significantly affects data quality, highlighting the importance of careful data selection for the instructional tuning of LLM. This section describes the scale and construction procedures of several commonly instruction tuning datasets.
@@
Self-instruct, created by Wang et al. (2023), which consists of 52,000 training and 252 testing instructions. Initial instructions from seed tasks were selected, categorized, and diversified using InstructGPT Ouyang et al. (2022b) to generate inputs and outputs through output-first or input-first strategies. Post-processing refined the dataset for uniqueness and relevance, providing a versatile resource for natural language processing applications.
Alpaca, created by Taori et al. (2023) which consists of 52,002 samples and is used to fine-tune LLaMA for the ability of instruction-following. Based on the techniques of Wang et al. (2023), the samples are generated by employing text-davinci-003.
@@
WizardLM, created by Xu et al. (2023a), which consists of 250,000 samples generated by the evolutionary algorithms. Two types of algorithms (depth evolution and breadth evolution) are used to enhance the complexity and scope of basic instructions, generating more complex and diverse high-quality instruction data by ChatGPT
@@
LIMA, created by Zhou et al. (2023a), which consists of 1,000 training samples, 300 test samples and 50 development samples. In addition to the manually authored samples, samples collected from Q&A websites are strictly selected by human. Despite its modest scale, LIMA stands out for its meticulous compilation and design. LLMs fine-tuned on LIMA show remarkable ability in following instructions and adapting to unseen tasks.
Dolly-v2, created by Conover et al. (2023) with 15,000 instructions, involving various tasks like brainstorming, classification, QA, and summarization. Employees author (prompt, response) pairs manually. They are restricted to using Wikipedia and advised against using web sources or generative AI for crafting response.
@@
P3, created by Sanh et al. (2022), which integrates 170 NLP datasets with 2,052 prompts. These prompts, also known as task templates, convert conventional NLP tasks like question answering or text classification into natural language input-output pairs. The P3 dataset itself is formed by randomly selecting prompts from PromptSource and organizing data into triplets of inputs, answer choices, and targets.
3 Data Selection methods
@@
Formally, define an instruction dataset $X$ with a size of $n$, where $X=\{x_{1},x_{2},\dots,x_{n}\}$, and each $x_{i}$ represents an instance of instruction fine-tuning data. To employ a specific instruction data selection method $\pi$ from $X$ and choose a subset $S_{\pi}^{(m)}$ with a size of $m$Liu et al. (2023), we then use a predefined evaluation metric $Q$ to assess the quality of $S_{\pi}^{(m)}$. The obtained subset quality, as measured by the evaluation metric, can gauge the effectiveness of the chosen instruction data selection method. The process of designing the selection method can be seen as:
$S^{(m)}_{\pi}=\pi\left(X\right)$ (1)
$\pi^{*}=\arg\max_{\pi}Q\left(S^{(m)}_{\pi}\right)$ (2)
@@
3.1 Methods based on System of indicators
Methods employing a metric set system directly identify multiple metrics $I_{1},I_{2},\dots,I_{n}$, thereby establishing a comprehensive set of metrics. Each metric within this set is defined by a specific calculation formula. Notably, certain metrics might leverage deep learning techniques for feature extraction from datasets, which are, in essence, metric forms. These metrics facilitate the computation of individual scores for data instances, denoted as $score_{ij}=I_{i}(x_{j})$. These scores are subsequently instrumental in developing a more robust metric set system.
$score_{j}=G\left(I_{1}(x_{j}),I_{2}(x_{j}),\dots,I_{n}(x_{j})\right)$ (3)
Once established, the metric set system can be directly utilized to compute scores for each data instance within the dataset $X$. By establishing suitable thresholds, this system facilitates the selective inclusion of data based on their respective scores:
$S_{\pi}=\{x|G(x)>\tau\}$ (4)
@@ 3.1 Methods based on System of indicators
Cao et al. (2023) introduce INSTRUCTMINING, a linear rule-based method for assessing the quality of instructional data. This approach initially identifies key natural language metrics, such as instruction length, perplexity, reward scores, KNN-iReimers and Gurevych (2019); Dong et al. (2011), among others. These metrics are subsequently utilized to develop a linear equation. To explore the correlation between data quality and these metrics, and to define the parameters of equation, comprehensive fine-tuning experiments are conducted. Diverse quality datasets are partitioned into subsets, combined, and then employed in the fine-tuning of large-scale models. Quality labels for each subset are derived by evaluating performance of the model on a test set. The least squares method is applied to these experimental outcomes to estimate the parameters in INSTRUCTMINING. This involves fitting a linear equation to the loss observed when evaluating the model against the test set.Once the parameters have been determined, this formula can be utilized to calculate the quality of instructions, thereby facilitating data selection.
@@
Wei et al. (2023) present InstructionGPT-4, a data selection method for multimodal large model fine-tuning. It surpasses MiniGPT-4Zhu et al. (2023) in various assessments using less data. The first step, metrics like the CLIP ScoreRadford et al. (2021), instruction length, and others are used. Visual and textual data are encoded into vectors and then subjected to dimensional reduction, which is regarded as special metrics. These metrics are combined into a vector $e$. The vector is then fed to a trainable data selector, such as a multi-layer perceptron or a self-attention network. This approach, similar to Cao et al. (2023) in calculating quality labels, employs clustering algorithms to segment datasets. Quality labels are assigned post fine-tuning and evaluation of each subset.
@@
Zhou et al. (2023b) introduce the DQ method, which is an innovative data compression technique for large-scale computer vision datasets, but it has been adapted for use in the LLM domain as well. The approach involves several key steps: Initially, a gain function $P(x)$ is defined,
$P(x)=\sum_{p\in S}||f(p)-f(x)||^{2}_{2}-\sum_{p\in D\setminus S}||f(p)-f(x)||^{2}_{2}$ (5)
@@
3.2 Methods based on Trainable LLMs
This section outlines the use of trainable LLMs, such as LLaMa, for developing computation formulas in data selection processes. The LLM, serving as a trainable data selector, processes and assigns scores to each piece of instruction fine-tuning data.
$score_{i}=LLM_{trainable}(x_{i})$ (6)
This method not only concentrates on analyzing individual instructions but also underscores the necessity of synchronizing data selection with the functionality of the large model employed for fine-tuning. Subsequent sections will detail specific methodologies.
Li et al. (2023a) propose the IFD method. It hypothesizes that LLMs can initially learn to recognize instructions from carefully selected data, enhancing their ability to assess broader dataset quality and estimate instruction-following difficulty. Initially, the method involves fine-tuning an LLM on a small, clustered subset of the instruction dataset to develop basic instruction-following skills. A novel metric, ”Instruction Following Difficulty (IFD),” is then introduced to assess the challenge of responding to specific instructions. IFD compares response quality without specific instructions. The Conditioned Answer Score is defined as:
$s_{\theta}(A|Q)=\frac{1}{N}\sum_{i=1}^{n}logP(w_{i}|Q,w_{1},\dots,w_{i-1})$ (7)
This score evaluates alignment of the model with the correct answer post-instruction (Q), considering the influence of generating the answer (A). The final IFD is:
$r_{\theta}(Q,A)=\frac{s_{\theta}(A|Q)}{s_{\theta}(A)}$ (8)
The IFD score quantifies challenges of the model for each sample. By setting an IFD score threshold, specific instructions are selected for initial LLM pre-training, resulting in a refined model.

### v2

@@ Abstract.
Instruction tuning is a vital step of training large language models (LLMs), so how to enhance the effect of instruction tuning has received increased attention. Existing works indicate that the quality of the dataset is more crucial than the quantity during instruction tuning of LLMs. Therefore, recently a lot of studies focus on exploring the methods of selecting high-quality subset from instruction datasets, aiming to reduce training costs and enhance the instruction-following capabilities of LLMs. This paper presents a comprehensive survey on data selection for LLM instruction tuning. Firstly, we introduce the wildly used instruction datasets. Then, we propose a new taxonomy of the data selection methods and provide a detailed introduction of recent advances, and the evaluation strategies and results of data selection methods are also elaborated in detail. Finally, we emphasize the open challenges and present new frontiers of this task.
copyright: ccdoi: 10.1613/jair.1.xxxxxjournalvolume: 4article: 6publicationmonth: 8journalyear: 2025
1. Introduction
Large language models (LLMs), such as PaLM (Chowdhery et al., 2023), GPT-4 (OpenAI, 2023), and LLaMA (Touvron et al., 2023), have demonstrated impressive capabilities across a wide range of natural language understanding and generation tasks. Central to their success is instruction tuning—a process that aligns model outputs with human intent by fine-tuning pretrained models on curated (instruction, response) pairs (Ouyang et al., 2022b; Zhang et al., 2023). This technique not only improves the controllability and safety of LLMs but also enables them to adapt to specific domains with minimal computational overhead.
In the existing works (Ghosh et al., 2024; Wei et al., 2021; Longpre et al., 2023), instruction tuning mainly focuses on constructing large-scale instruction-following datasets. These works can be categorized into two types: (1) converting annotated NLP datasets into instruction-style formats using templates (e.g., P3 (Sanh et al., 2022) and FLAN (Wei et al., 2021)), and (2) automatically generating instruction-response pairs using powerful models like GPT-3.5 (e.g., Self-Instruct (Wang et al., 2023)). While these works have
@@ 1. Introduction
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=12em,font=,, where level=2text width=12em,font=,, where level=3text width=8em,font=,, where level=4text width=5em,font=,, [ Data Selection for Instruction Tuning, ver [Instruction Sets (§2) [ Alpaca (Taori et al., 2023), WizardLM (Xu et al., 2023a), Dolly-V2 (Conover et al., 2023), P3 (Sanh et al., 2022), LIMA (Zhou et al., 2023a), Self-Instruct (Wang et al., 2023) , leaf, text width=36em ] ] [Data Selection Methods (§3) [System of indicators (§3.1) [ INSTRUCTMINING (Cao et al., 2023), DQ (Zhou et al., 2023b),
@@
InstructionGPT-4 (Wei et al., 2023) , leaf, text width=18em ] ] [Trainable LLMs (§3.2) [ IFD (Li et al., 2024c), Instruction Backtranslation (Li et al., 2024b),
Nuggets (Li et al., 2024a), DIVERSEEVOL (Wu et al., 2023),
@@
TEGIT (Chen et al., 2023a), ActiveIT (Kung et al., 2023) , leaf, text width=18em ] ] [Powerful LLMs (§3.3) [ AlpaGasus (Chen et al., 2024), INSTAG (Lu et al., 2024),
LIFT (Xu et al., 2023b), DEITA (Liu et al., 2024),
@@
tree-instruct (Zhao et al., 2023), WaveCoder (Yu et al., 2023) , leaf, text width=18em ] ] [Small Models (§3.4) [ MoDS (Du et al., 2023), Coreset-based Selection (Chen et al., 2023b) , leaf, text width=18em ] ] ] [Evaluation Methods (§4) [ Wining rate (§4.1), Inner Comparison (§4.2), External Comparison (§4.3), leaf, text width=36em ] ] ] ]
Figure 1. Overview of Data Selection for LLM Instruction Tuning.
greatly expanded the availability of instruction data, they often suffer from issues related to data redundancy, limited diversity, and inconsistent quality. Moreover, these works primarily rely on a large volume of data, which can lead to a significant waste of computational resources and electricity.
In contrast to the previous approaches primarily focused on large-scale data construction (Taori et al., 2023; Xu et al., 2023a; Conover et al., 2023), LIMA (Zhou et al., 2023a) proposes an approach centered on careful data selection. Their findings demonstrated that a model fine-tuned on merely 1,000 carefully curated instruction examples could achieve performance comparable, and in some cases superior, to models trained on vastly larger, often automatically generated, datasets. This seminal work underscored a critical insight: the quality of instruction data can significantly outweigh its quantity in determining the efficacy of instruction tuning. Consequently, this discovery has prompted a shift in research focus, moving from the large-scale generation of instruction data towards developing better methods for constructing and selecting high-quality, impactful, and efficient datasets.
In these data selection approaches, manual instruction data selection often involves high costs and introduces human bias. As a result, automated methods are crucial for instruction data selection. These methods have shown promising results. For instance, the IFD(Li et al., 2024c) method significantly outperforms the Alpaca model by using only about 5% of the Alpaca dataset, and it also surpasses the WizardLM model by about 10%. Using high-quality subsets for fine-tuning not only enhances the instruction-following skills of LLMs but also significantly reduces computational costs and time.
However, the data selection task for LLM instruction tuning is challenging due to the complex factors and multidimensional considerations involved. For example, it is difficult to assess the quality of individual instructions and ensure the overall diversity of the selected data. Another challenge is reducing costs and improving the efficiency of the selection process. In light of these factors, various data selection methods have been developed. Some methods use a system of indicators to evaluate individual data points, while others rely on trainable LLMs or powerful external LLMs. These exploit the LLMs’ own capabilities to select instructions. Additionally, methods that use smaller models and design a comprehensive process to achieve a balanced effect across all aspects are noteworthy.
Despite these promising advances, the field remains fragmented, with no unified understanding of the principles, methods, and trade-offs involved in instruction data selection. In this survey, we aim to fill this gap by providing a systematic overview of instruction data selection for LLMs. We categorize existing methods, analyze their core ideas and assumptions, and highlight open challenges such as the lack of standardized evaluation protocols, inefficiencies in processing large-scale instruction data, and the heavy dependence on powerful foundation models for selection. By synthesizing recent progress and identifying future directions, we hope to establish a foundation for further research and practical development in instruction tuning.
@@
As illustrated by Fig. 1, this paper provides a comprehensive review of existing works on data selection methods for the instruction tuning of LLM. To facilities the community, we maintain a paper list [^1] https://github.com/Bolin97/awesome-instruction-selector, collect commonly instruction sets for data selection. Section 2 describes the mainstream datasets with different source and construction methods used in instruction tuning and Section 3 describes four types data selection methods in details: indicators sets, trainable LLMs, powerful LLMs and small models. Section 4 presents the evaluation methods and shows the results of different instruction selection methods. Section 5 summarizes the paper and emphasize the open challenges and future direction of the instruction selection.
2. Instruction Datasets
@@
Various instruction tuning datasets (e.g. Self-Instruct and Alpaca), generated by LLMs, offer a wealth of samples without human labor, but their data quality depends on the performance of LLMs and is uncertain. Conversely, the manually curated datasets (e.g. LIMA and Dolly) obtain higher quality through meticulous human selection but are potentially influenced by human biases. Alternative dataset construction methods, like prompt mapping and evol-instruct, aim to enhance dataset quality and diversity but introduce new challenges in quality assurance. This variability in dataset construction and sourcing significantly affects data quality, highlighting the importance of careful data selection for the instructional tuning of LLM. This section describes the scale and construction procedures of several commonly instruction tuning datasets:
@@
Self-instruct (Wang et al., 2023) consists of 52,000 training and 252 testing instructions. Initial instructions from seed tasks were selected, categorized, and diversified using InstructGPT (Ouyang et al., 2022a) to generate inputs and outputs through output-first or input-first strategies. Post-processing refined the dataset for uniqueness and relevance, providing a versatile resource for natural language processing applications.
Alpaca (Taori et al., 2023) is a dataset of 52,002 instruction-response pairs generated using the Self-Instruct framework (Wang et al., 2023) with the text-davinci-003 model. It is designed to fine-tune LLaMA for improved instruction-following capabilities.
@@
WizardLM (Xu et al., 2023a) is a dataset containing 250,000 instruction-response pairs generated via evolutionary algorithms. Two types of algorithms (depth evolution and breadth evolution) are used to enhance the complexity and scope of basic instructions, generating more complex and diverse high-quality instruction data by ChatGPT.
@@
LIMA (Zhou et al., 2023a) is a dataset consisting of 1,000 training samples, 300 test samples, and 50 development samples. In addition to the manually authored samples, samples collected from Q&A websites are strictly selected by human. Despite its modest scale, LIMA stands out for its meticulous compilation and design. LLMs fine-tuned on LIMA show remarkable ability in following instructions and adapting to unseen tasks.
DollyV2 (Conover et al., 2023) is a dataset of 15,000 human-authored instruction-response pairs covering tasks such as brainstorming, classification, question answering, and summarization. The data was created by employees using only Wikipedia as a source, with restrictions against using web content or generative AI tools when crafting responses.
@@
P3 (Sanh et al., 2022) is a large-scale prompt collection that integrates 170 NLP datasets with 2,052 handcrafted prompts, enabling unified training and evaluation across diverse language tasks. These prompts, also known as task templates, convert conventional NLP tasks like question answering or text classification into natural language input-output pairs. The P3 dataset itself is formed by randomly selecting prompts from PromptSource and organizing data into triplets of inputs, answer choices, and targets.
3. Data Selection methods
@@
Formally, define an instruction dataset $X$ with a size of $n$, where $X=\{x_{1},x_{2},\dots,x_{n}\}$, and each $x_{i}$ represents an instance of instruction fine-tuning data. To employ a specific instruction data selection method $\pi$ from $X$ and choose a subset $S_{\pi}^{(m)}$ with a size of $m$ , we then use a predefined evaluation metric $Q$ to assess the quality of $S_{\pi}^{(m)}$. The obtained subset quality, as measured by the evaluation metric, can gauge the effectiveness of the chosen instruction data selection method. The process of designing the selection method can be seen as:
(1) $S^{(m)}_{\pi}=\pi\left(X\right)$
(2) $\pi^{*}=\arg\max_{\pi}Q\left(S^{(m)}_{\pi}\right)$
@@
3.1. Methods Based on a System of Indicators
Methods based on a system of indicators aim to define a comprehensive set of evaluation metrics $I_{1},I_{2},\dots,I_{n}$, each governed by a specific computational formula. These metrics are designed to assess different aspects of data quality or relevance. In certain cases, the metrics may incorporate deep learning models for feature extraction, which function as learned indicators.
For a given data instance $x_{j}$, each metric $I_{i}$ produces a score $\text{score}_{ij}=I_{i}(x_{j})$. These individual metric scores are then aggregated using a combination function $G(\cdot)$, resulting in an overall score:
(3) $\text{score}_{j}=G\left(I_{1}(x_{j}),I_{2}(x_{j}),\dots,I_{n}(x_{j})\right)$
Once the scoring function $G$ is defined, the system can be applied to the entire dataset $X$ to evaluate each data instance. Based on a predefined threshold $\tau$, a high-quality subset $S_{\pi}$ can be selected as follows:
(4) $S_{\pi}=\{x\in X\mid G(x)>\tau\}$
This indicator-based framework provides a structured and scalable approach to data selection, enabling the integration of both heuristic and learned metrics.
@@ 3.1. Methods Based on a System of Indicators
INSTRUCTMINING (Cao et al., 2023) is a linear rule-based method for assessing the quality of instructional data. This approach initially identifies key natural language metrics, such as instruction length, perplexity, reward scores, KNN-i(Reimers and Gurevych, 2019; Dong et al., 2011), among others. These metrics are subsequently utilized to develop a linear equation. To explore the correlation between data quality and these metrics, and to define the parameters of equation, comprehensive fine-tuning experiments are conducted. Diverse quality datasets are partitioned into subsets, combined, and then employed in the fine-tuning of large-scale models. Quality labels for each subset are derived by evaluating performance of the model on a test set. The least squares method is applied to these experimental outcomes to estimate the parameters in INSTRUCTMINING. This involves fitting a linear equation to the loss observed when evaluating the model against the test set.Once the parameters have been determined, this formula can be utilized to calculate the quality of instructions, thereby facilitating data selection.
@@
InstructionGPT-4 (Wei et al., 2023) is a data selection method designed for fine-tuning multimodal large language models. It surpasses MiniGPT-4 (Zhu et al., 2024) in various assessments using less data. The first step, metrics like the CLIP Score(Radford et al., 2021), instruction length, and others are used. Visual and textual data are encoded into vectors and then subjected to dimensional reduction, which is regarded as special metrics. These metrics are combined into a vector $e$. The vector is then fed to a trainable data selector, such as a multi-layer perceptron or a self-attention network. This approach, similar to the method (Cao et al., 2023) of calculating quality labels, employs clustering algorithms to segment datasets. Quality labels are assigned post fine-tuning and evaluation of each subset.
@@
DQ (Zhou et al., 2023b) is an innovative data compression approach originally developed for large-scale computer vision datasets, and later adapted for use in the LLM domain. The approach involves several key steps: Initially, a gain function $P(x)$ is defined as:
(5) $P(x)=\sum_{p\in S}||f(p)-f(x)||^{2}_{2}-\sum_{p\in D\setminus S}||f(p)-f(x)||^{2}_{2},$
@@
3.2. Methods Based on Trainable LLMs
This section introduces methods that leverage trainable large language models (LLMs) to develop scoring functions for instruction data selection. In these methods, the LLM itself is fine-tuned or adapted to act as a data selector, assigning a score to each instruction instance:
(6) $\text{score}_{i}=\text{LLM}_{\text{trainable}}(x_{i})$
Unlike indicator-based methods, which rely on predefined or hand-crafted metrics, trainable LLM-based methods learn to evaluate instruction quality directly from data. These methods not only focus on scoring individual instructions but also emphasize the alignment between the data selection process and the downstream performance of the fine-tuned model. Specific methods and frameworks that implement this paradigm are discussed in the following sections.
Instruction following difficulty (IFD) (Li et al., 2024c) represents a trainable data selection approach grounded in the hypothesis that large language models (LLMs) can first be trained on a carefully curated subset of instruction data to acquire basic instruction-following capabilities. This foundational skill then enables the model to better estimate instruction difficulty and overall data quality.
The method begins by fine-tuning an LLM on a small, clustered subset of the instruction dataset. It then introduces a novel metric, also named IFD, which quantifies the challenge of responding to a given instruction. Specifically, it compares the model’s ability to generate the correct answer both with and without the given instruction. The Conditioned Answer Score is computed as:
(7) $s_{\theta}(A|Q)=\frac{1}{N}\sum_{i=1}^{n}\log P(w_{i}\mid Q,w_{1},\dots,w_{i-1}),$
where $Q$ is the instruction and $A$ is the target answer consisting of tokens $w_{1},\dots,w_{n}$. This measures the model’s likelihood of generating $A$ when conditioned on $Q$.
The final IFD score is defined as the ratio between the conditioned and unconditioned scores:
(8) $r_{\theta}(Q,A)=\frac{s_{\theta}(A|Q)}{s_{\theta}(A)},$
where $s_{\theta}(A)$ is the score of generating $A$ without conditioning on $Q$. Higher IFD scores indicate samples where the instruction significantly influences the output, thereby reflecting greater instruction-following difficulty. By applying a threshold on the IFD score, the method selects informative instructions to further refine the model in subsequent fine-tuning stages.